OpenStudy (anonymous):

Simplify the expression: (2x-3y)^2/(3y-2x)^2

OpenStudy (anonymous):

\[\frac{4 x^2-12 x y+9 y^2}{4 x^2-12 x y+9 y^2}=1 \]

OpenStudy (anonymous):

how did you get this

OpenStudy (anonymous):

\[\text{Expand}\left[(2 x-3 y)^2\right]=4 x^2-12 x y+9 y^2 \]\[\text{Expand}\left[(3 y-2 x)^2\right]=4 x^2-12 x y+9 y^2 \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

you are welcomed